OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -7, an = 2 • an-1

Nnesha (nnesha):

we did it yesterday right ?

OpenStudy (anonymous):

It's not the same problem. I'm just confused with these types.

Nnesha (nnesha):

same method \[\huge\rm a_n=2 \times a_{n-1}\] n= term that you need a_1 first term

Nnesha (nnesha):

to find 2nd term replace n by 2 \[\huge\rm a_2=2 \times a_{2-1}\]

OpenStudy (anonymous):

Ok...so would this be right? -7, -14, -28, -56, -112, -224

Nnesha (nnesha):

yep